# Reflectionless and echo modes in asymmetric Damour-Solodukhin wormholes

### v2

@@ V.2 The waveforms of RSMs and echo modes
where $\sigma_{0}=1$, $y_{0}=-\frac{1}{4}$, and $\omega_{0}=-4+2i$. This represents a pulse localized at $y=y_{0}$, with a spread in the frequency domain. Physically, this source term encodes the information about the initial/boundary conditions from the Fourier [90] and Laplace [86, 8] transform perspectives. As detailed further in Appx. E, the numerator $\omega^{2}$ suppresses the pole at the origin, the denominator ensures faster convergence of the inverse Fourier transform, and the additional pole at $\omega=\omega_{0}$ plays no role since the contour integration follows a large circle in the lower half of the complex frequency plane. The resulting frequency-domain waveform is given by
@@ VI Concluding remarks
We argue that such an analysis for asymmetric wormholes is pertinent, since realistic wormholes, if they exist at all, are unlikely to be symmetric, and poles lying closer to the real frequency axis exert a more significant impact on time-domain waveforms. Recent developments on gravitational wave echoes, spectral instability, echo modes, and Regge poles are closely related to, and partly motivated by, the ongoing efforts in black hole spectroscopy. Given the arguments that greybody factors may provide more relevant observables in light of black hole spectral instability, the notion of RSMs was proposed and explored. The present study elaborates further on this concept, offering an explicit comparison with echo modes. It is understood that echo modes and RSMs are two perspectives that complement one another and together provide effective tools for describing the underlying phenomenon.

### v3

@@ V.2 The waveforms of RSMs and echo modes
where $\sigma_{0}=1$, $y_{0}=-\frac{1}{4}$, and $\omega_{0}=-4+2i$. This represents a pulse localized at $y=y_{0}$, with a spread in the frequency domain. Physically, this source term encodes the information about the initial/boundary conditions from the Fourier [95] and Laplace [91, 7] transform perspectives. We also note that the specific form of Eq. (78) is not essential. As detailed further in Appx. E, the numerator $\omega^{2}$ suppresses the pole at the origin, the denominator ensures faster convergence of the inverse Fourier transform, and the additional pole at $\omega=\omega_{0}$ plays no role since the contour integration follows a large circle in the lower half of the complex frequency plane. The resulting frequency-domain waveform is given by
@@ VI Concluding remarks
We argue that such an analysis for asymmetric wormholes is pertinent, since realistic wormholes, if they exist at all, are unlikely to be symmetric, and poles lying closer to the real frequency axis exert a more significant impact on time-domain waveforms. Recent developments on gravitational wave echoes, spectral instability, echo modes, and Regge poles are closely related to, and partly motivated by, the ongoing efforts in black hole spectroscopy. Given the arguments that greybody factors may provide more relevant observables in light of black hole spectral instability, the notion of RSMs was proposed and explored. The present study further elaborates on this concept, primarily in the context of Damour-Solodukhin wormholes, and offers an explicit comparison with echo modes. It is worth noting that recently in [88], the RSMs, referred to by the authors as total absorption or virtual absorption modes, were investigated for the Regge–Wheeler effective potential, for which the algebraically special mode is a well-known solution, and also in higher dimensions. It is understood that echo modes and RSMs are two perspectives that complement one another and together provide effective tools for describing the underlying phenomenon.